Consider C^n, with the standard inner product. Find an orthonormal basis for the subspace spanned by v1=(1, 0, i) and v2=(2, 1, 1+i)
1
Expert's answer
2020-12-20T18:25:29-0500


Orthonormalize the set of the vectors v1 = "\begin{bmatrix}
 1 \\
 0 \\
i
\end{bmatrix}" and v2 = "\begin{bmatrix}
 2 \\
 1 \\ 
1 + i
\end{bmatrix}"

According to the Gram-Schmidt process, uk = vk - "\sum_{j - 1}^{k - 1} proj_{uj}(v_k)" where

"proj_u(v) = \frac{u*v}{u * u}u"

The normalized vector is "e_k = \frac{u_k}{\sqrt{u_k*u_k}}"


step1: "u_1 = v_! =" "\begin{bmatrix}
 1 \\
 0 \\
i
\end{bmatrix}"


"e_1 = \frac{u_1}{\sqrt{u_1*u_1}} = \begin{bmatrix}
 \frac{\sqrt{2}}{2} \\
 0 \\
 \frac{\sqrt{2}i}{2} 
\end{bmatrix}"



step2: "u_2 = v_2 - \frac{u_1*v_2}{u_1*u_1}u_1 = \begin{bmatrix}
 \frac{1}{2} - \frac{i}{2} \\
 1 \\
\frac{3}{2} - \frac{i}{2} 
\end{bmatrix}"


"e_2 = \frac{u_2}{\sqrt{u_2*u_2}} = u_2 = v_2 - \frac{u_1*v_2}{u_1*u_1}u_1 = \begin{bmatrix}
 \frac{1}{4} - \frac{i}{4} \\
 \frac{1}{2} \\
\frac{3}{4} - \frac{i}{4} 
\end{bmatrix}"


Answer: the set of orthonormal vectors is


"e_1 = \begin{bmatrix}
 \frac{\sqrt{2}}{2} \\
 0 \\
 \frac{\sqrt{2}i}{2} 
\end{bmatrix}", "e_2 = \begin{bmatrix}
 \frac{1}{4} - \frac{i}{4} \\
 \frac{1}{2} \\
\frac{3}{4} - \frac{i}{4} 
\end{bmatrix}"